OpenStudy (anonymous):

Solve the following system of equations by graphing. x + 2y = 4 2x + 4y = 8 What is the solution? (4,2) no solution infinite solutions none of the above

OpenStudy (anonymous):

so its none of those

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry would be (4,2)

jimthompson5910 (jim_thompson5910):

notice how everything in the 2nd equation is double of the numbers in the first equation

OpenStudy (anonymous):

its not (4,2) cause i had that before and i got it wrong

OpenStudy (anonymous):

if you plug in 0 for x and then for y then you get 4,2

OpenStudy (anonymous):

but this is supposed to be solved by graphing..

OpenStudy (anonymous):

and i did get (4,2) and it is wrong

jimthompson5910 (jim_thompson5910):

(4,2) is a solution, but it's not the only solution

OpenStudy (anonymous):

so it should of been infinate solutions

jimthompson5910 (jim_thompson5910):

yes, the reason being because the two equations are really the same equation

jimthompson5910 (jim_thompson5910):

the second is just the first (but everything is doubled)

OpenStudy (anonymous):

how so

OpenStudy (anonymous):

oh gotcha

jimthompson5910 (jim_thompson5910):

x + 2y = 4 2*x + 2*2y = 2*4 2x + 4y = 8

OpenStudy (anonymous):

gotcha (:

OpenStudy (anonymous):

so hat other soultions will u get

OpenStudy (anonymous):

what ?

jimthompson5910 (jim_thompson5910):

anything that satisfies x+2y = 4

jimthompson5910 (jim_thompson5910):

so (0,2) is a solution x+2y = 4 0+2*2 = 4 4 = 4 ------------------------ (-2,3) is another solution x+2y = 4 -2+2(3) = 4 -2+6 = 4 4 = 4

OpenStudy (anonymous):

alright thanks (:

jimthompson5910 (jim_thompson5910):

yw